### v1

Abstract
In this work, we consider constrained stochastic optimization problems under hidden convexity, i.e., those that admit a convex reformulation via non-linear (but invertible) map $c(\cdot)$. A number of non-convex problems ranging from optimal control, revenue and inventory management, to convex reinforcement learning all admit such a hidden convex structure. Unfortunately, in the majority of applications considered, the map $c(\cdot)$ is unavailable or implicit; therefore, directly solving the convex reformulation is not possible. On the other hand, the stochastic gradients with respect to the original variable are often easy to obtain. Motivated by these observations, we examine the basic projected stochastic (sub-) gradient methods for solving such problems under hidden convexity. We provide the first sample complexity guarantees for global convergence in smooth and non-smooth settings. Additionally, in the smooth setting, we improve our results to the last iterate convergence in terms of function value gap using the momentum variant of projected stochastic gradient descent.
1 Introduction
We study constrained stochastic optimization
$\min_{x\in\mathcal{X}}~{}F(x):={\mathbb{E}}_{\xi\sim\mathcal{D}}\left[f(x,\xi)\right],$ (1)
where $\mathcal{X}$ is a closed convex subset of $\mathbb{R}^{d}$, $\xi$ is a random variable following an unknown distribution $\mathcal{D}$, and $F(\cdot)$ is possibly non-convex in $x$. Our central structural assumption about (1) is that it admits a convex reformulation of the form
$\min_{u\in\mathcal{U}}~{}H(u):=F(c^{-1}(u)),$ (2)
where $H(\cdot)$ is a convex function defined on a closed convex set $\mathcal{U}\subset\mathbb{R}^{d}$, and $c:\mathcal{X}\rightarrow\mathcal{U}$ is an invertible map (with its inverse denoted by $c^{-1}(\cdot)$). This property is often referred to as hidden convexity and frequently appears in various modern applications, for example, policy optimization in convex reinforcement learning and optimal control [92, 80], generative models [50], supply chain and revenue management [36, 15], training neural networks [83, 30]. In the optimization literature, hidden convexity has been identified much earlier and dates back to at least 1990s in the context of quadratic optimization [78, 8]. Since then, several works have developed various tools to identify such property [7, 6]. More recently, hidden convexity has been established for a wider classes of non-convex programs [81, 84, 14] and for non-monotone games [82, 62], to name just a few. Despite the existence of the convex reformulation, the transformation function $c(\cdot)$ is usually hard to compute or even unknown, and one cannot readily solve the convex reformulation. This motivates the use of (sub-) gradient methods that optimize $F(\cdot)$ directly over the variable $x\in\mathcal{X}$. Perhaps, the most basic algorithm is the projected stochastic (sub-) gradient method (SM). Starting from $x^{0}\in\mathcal{X}$, SM generates a sequence $x^{t}$ via
$x^{t+1}=\Pi_{\mathcal{X}}(x^{t}-\eta g(x^{t},\xi^{t})),$ (3)
where $g(x^{t},\xi^{t})$ denotes an unbiased estimate of the (sub-) gradient of $F(\cdot)$ at a point $x^{t}$, and $\Pi_{\mathcal{X}}(\cdot)$ is the Euclidean projection onto a convex set $\mathcal{X}$, see Section 1.2 for details. When $F(\cdot)$ is differentiable, this reduces to Projected SGD. Stochastic (sub-) gradient methods and their numerous variants have a long history of development since the first works on stochastic approximation appeared in 1950s [73, 49, 9, 18]. Their analyses are richly documented for addressing convex problems and general nonconvex problems (refer to Appendix C for detailed summary); however, their convergence behaviors when dealing with hidden convexity still elude precise understanding.
Although hidden convexity has been previously identified in certain applications, the analysis of gradient methods under this condition is mostly done on a case by case basis for specific applications and often requires strong additional assumptions [93, 4, 16, 15]. In this work, we formally consider general-purpose stochastic optimization under hidden convexity and study the sample complexities for solving such problems through projected stochastic (sub-) gradient method and the like.
Contributions:
  1. 1.
    We identify key properties of hidden convex optimization and demonstrate how these conditions can be used to derive global convergence of gradient methods.
  2. 2.
    In the general non-differentiable case, we analyze convergence of the projected stochastic sub-gradient method (SM) and obtain $\widetilde{\mathcal{O}}\left(\frac{\ell}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{\ell G_{F}^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right)$ sample complexity for driving the Moreau envelope of (1) $\varepsilon$-close to the optimal value in expectation. Here $\ell$ is weak convexity parameter of $F(\cdot)$, $G_{F}$ is the bound on (stochastic) subgradients, and $\mu_{c}$ is the modulus of hidden convexity of $F(\cdot)$ that relates to the invertibility of mapping $c(\cdot)$; see Section 2 and Section 4 for details. To our knowledge, it is the first result to address the non-differentiable setting under hidden convexity.
  3. 3.
    Next, we specialize our results to the differentiable smooth setting, and obtain a similar sample complexity for Projected SGD, replacing $G_{F}$ by the variance of stochastic gradients $\sigma$. Furthermore, we analyze the momentum variant of Projected SGD improving our result to the last iterate convergence in terms of the function value gap, i.e., we have ${\mathbb{E}}\left[F(x^{T})-F^{*}\right]\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{L}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{L\sigma^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right)$ iterations/stochastic gradient calls, where $L$ is the Lipschitz constant of $\nabla F(x)$.
  4. 4.
    In the presence of strong convexity of the reformulated problem, we further improve the sample complexity for all above mentioned algorithms. For instance, Projected SGD attains $\widetilde{\mathcal{O}}\left(\frac{L}{\mu_{c}^{2}\mu_{H}}+\frac{L\sigma^{2}}{\mu_{c}^{4}\mu_{H}^{2}}\frac{1}{\varepsilon}\right)$ sample complexity for achieving an $\varepsilon$-optimal solution, where $\mu_{H}>0$ corresponds to the strong convexity of $H(\cdot)$.
Importantly, we show that all studied algorithms provably converge in online fashion, using only one stochastic gradient at every iteration.
1.1 Related work
The most closely related works are [92, 93, 4, 16, 40], which analyze gradient methods under similar structural assumptions in the context of specific applications.
Policy gradient methods in RL. Several work [92, 93, 4] exploited properties similar to hidden convexity in reinforcement learning (RL) applications and analyzed policy gradient (PG) type methods with global convergence guarantees. In [92], the authors considered a PG method with projection, but it is only limited to the case where the exact gradients are available. It is unclear how to extend the technique in their work to the case of stochastic gradients with bounded variance (without resorting to large batches). Next, [93] considered the stochastic setting and proposed a variance-reduced PG method with truncation using large batches of trajectories. Recently, [4] removed the requirement for large batches using a normalized variance-reduced PG method. However, their results are difficult to extend to the constrained case due to the normalization. Moreover, both works [93, 4] utilize variance reduced estimators, which require additional smoothness assumptions for theoretical analysis.
Stochastic gradient methods in revenue management. A different line of works [15, 16] considered hidden convex objectives in revenue management and studied global convergence of gradient-based methods over $\mathcal{X}$. For a special revenue management problem, [15] introduced a preconditioned gradient-based method that obtains an $\widetilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity under the assumptions that the domain $\mathcal{X}$ is a box constraint, the transformation function $c(x)={\mathbb{E}}\left[c(x,\xi)\right]$ is separable and the additional access to $c(x,\xi)$ is available. Leveraging the box constraint structure, [15] also analyzed Projected SGD and derive $\widetilde{\mathcal{O}}(\epsilon^{-4})$ sample complexity. In contrast, we show that Projected SGD can achieve a better $\widetilde{\mathcal{O}}(\epsilon^{-3})$ sample complexity for a general convex compact constraint $\mathcal{X}$, and further extend the results to non-smooth setting.
Nonconvex online learning. Recently, [40] considered a structural property similar to hidden convexity and imposed strong assumptions on the reparameterization map $c(\cdot)$ (see Assumptions 1, 2 and 4 therein) under which non-convex online gradient descent in the original space $\mathcal{X}$ is equivalent to online mirror descent for the (convex) reformulated problem. Such equivalence allows them to demonstrate an ${\mathcal{O}}(T^{\nicefrac{{2}}{{3}}})$ regret bound. Instead, we directly derive the last iterate convergence in the function value using a different technique and make less restrictive assumptions on $c(\cdot)$, which allows us to cover a wide range of applications.
Related structural assumptions. We mention that several other non-convex structural assumptions have been explored in optimization literature that also ensures global convergence, including essential strong convexity [56], quasar (strong) convexity [42], restricted secant inequality [90], error bounds [59], quadratic growth [10], Polyak-Łojasiewicz (PŁ) condition [71, 58] – also known as global Kurdyka-Łojasiweicz (KŁ) and gradient domination condition. For an in-depth discussion on the relationships between these properties, refer to  [47, 72] and references therein. Notably, the PŁ condition along with its various generalizations to constrained minimization such as Proximal-PŁ [47] and variational gradient dominance [85] have gained popularity in the recent years. The convergence of gradient-based methods under the PŁ-type conditions has been extensively analyzed, e.g., in the deterministic setting [47, 89] and the stochastic setting [32, 37, 48, 75, 54, 17, 88, 31]. Despite a few examples [35, 33, 26, 85] that show some variants of the PŁ condition hold, how to verify PŁ-like conditions for non-convex problems remains a big question in general. The situation becomes even more challenging, when dealing with constrained optimization and/or non-differentiable objectives, where a suitable generalization of the gradient dominance needs to be introduced and carefully studied. Unlike the PŁ condition, hidden convexity considered in this work is a very natural property, which easily extends to constrained optimization and non-differentiable objectives.
1.2 Notations and organization
[…]
A function $F:\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ is $\ell$-weakly convex ($\ell$-WC) if for any fixed $y\in\mathcal{X}$, $F_{\ell}(x,y):=F(x)+\frac{\ell}{2}\,\left\|x-y\right\|^{2}$ is convex in $x\in\mathcal{X}$. If $x\notin\mathcal{X}$, we assign $F(x)=+\infty$. The (Fréchet) sub-differential of $F$ at $x\in\mathcal{X}$ is $\partial F(x):=\left\{g\in\mathbb{R}^{d}\mid F(y)\geq F(x)+\langle g,y-x\rangle+o(\|y-x\|),\forall y\in\mathbb{R}^{d}\right\}.$ The elements $g\in\partial F(x)$ are called sub-gradients of $F$ at $x$, see [24] for alternative equivalent definitions of the sub-differential set for $\ell$-WC functions. A differentiable function $F:\mathcal{X}\rightarrow\mathbb{R}$ is $L$-smooth on $\mathcal{X}\subset\mathbb{R}^{d}$ if its gradient is $L$-Lipschitz continuous on the set $\mathcal{X}$, i.e., it holds $\left\|\nabla F(x)-\nabla F(y)\right\|\leq L\left\|x-y\right\|$ for all $x,y\in\mathcal{X}$. For a convex set $\mathcal{X}\subset\mathbb{R}^{d}$, the projection of a point $y\in\mathbb{R}^{d}$ onto $\mathcal{X}$ is $\Pi_{\mathcal{X}}(y):=\operatorname*{arg\,min}_{x\in\mathcal{X}}\left\|y-x\right\|$. We denote $\delta_{\mathcal{X}}$ as the indicator function of a set $\mathcal{X}\subset\mathbb{R}^{d}$ and define $\delta_{\mathcal{X}}(x)=0$ if $x\in\mathcal{X}$ and $\delta_{\mathcal{X}}(x)=+\infty$ otherwise. We define by $\mathcal{X}^{*}\subset\mathcal{X}$ the set of optimal points of $\min_{x\in\mathcal{X}}F(x)$ and by $F^{*}$ its optimal value. A point $\bar{x}\in\mathcal{X}$ is called a stationary point of a weakly convex function $F:\mathcal{X}\rightarrow\mathbb{R}$ if $0\in\partial(F+\delta_{\mathcal{X}})(\bar{x})$. For any function $\Phi$ and a real $\rho>0$, we define the Moreau envelope and the proximal mapping as follows
$\Phi_{1/\rho}(x):=\min_{y\in\mathbb{R}^{d}}\left\{\Phi(y)+\frac{\rho}{2}\left\|y-x\right\|^{2}\right\},\quad\operatorname{prox}_{\Phi/\rho}(x):=\operatorname*{arg\,min}_{y\in\mathbb{R}^{d}}\left\{\Phi(y)+\frac{\rho}{2}\left\|y-x\right\|^{2}\right\}.$
The rest of the paper is organized as follows. We formally introduce the hidden convex function class in Section 2 followed by motivating examples. The properties of hidden convex optimization, which are useful to analyze global convergence of gradient methods are in Section 3. Our main global convergence results are in Sections 4, 5, and 6 for subgradient methods, Projected SGD, and Projected SGD with momentum, respectively. The conclusion follows in Section 7.
2 Hidden Convex Problem Class
The existence of a convex reformulation (2) for the problem (1) signifies its representation as a compositional optimization in the form:
$\min_{x\in\mathcal{X}}~{}F(x):=H(c(x)).$ (4)
Formally, we make the following definition.
Definition 1.
The above problem is called hidden convex with modulus $\mu_{c}>0,\mu_{H}\geq 0$ (or function $F$ is hidden convex on $\mathcal{X}$) if its components satisfy the following underlying conditions.
  1. C.1
    The domain $\mathcal{U}=c(\mathcal{X})$ is convex, the function $H:\mathcal{U}\rightarrow\mathbb{R}$ satisfies that for all $u,v\in\mathcal{U}$ and $\lambda\in[0,1]$,
    $H((1-\lambda)u+\lambda v)\leq(1-\lambda)H(u)+\lambda H(v)-\frac{(1-\lambda)\lambda\mu_{H}}{2}\left\|u-v\right\|^{2},$ (5)
    and (2) admits a solution $u^{*}\in\mathcal{U}$.
  2. C.2
    The map $c:\mathcal{X}\rightarrow\mathcal{U}$ is invertible. There exists $\mu_{c}>0$ such that for all $x,y\in\mathcal{X}$ it holds
    $\left\|c(x)-c(y)\right\|\geq\mu_{c}\left\|x-y\right\|.$ (6)
In particular, if $\mu_{H}>0$, we say the above problem is $(\mu_{c},\mu_{H})$-hidden strongly convex.
Notice that the hidden convex problem class includes the convex problem as a special case when the transformation map $c(\cdot)$ is identical. In addition, it also includes many non-convex problems. For a simple example, let $0<\delta\leq 1$ and consider $\mathcal{X}=[\delta,1]$, $c(x)=x^{2}$, $H(u)=-u$. Then $F(x)=-x^{2}$, albeit concave, is hidden convex on $\mathcal{X}$ by the construction. Another simple example considers $0<\delta<\pi$ and $\mathcal{X}=[\delta,2\pi-\delta]$, $c(x)=cos(x)$, $H(u)=u$. The obtained composition $F(x)=cos(x)$ is also hidden convex on $\mathcal{X}$, although it is both non-convex and non-concave on $\mathcal{X}$.
In what follows, we present several more practical problems, which belong to our hidden convex class.
2.1 Non-linear least squares [67, 66, 28]
Consider solving a system of nonlinear equations under a box constraint, e.g., $c(x)=0$ with $c(x)=(c_{1}(x),\ldots,c_{d}(x))^{\top}$ for $x\in\mathcal{X}=[0,D]^{d}$, $D>0$. This problem can be equivalently formulated as
$\min_{x\in[0,D]^{d}}\sum_{i=1}^{d}c_{i}^{2}(x)\qquad\text{or }\qquad\min_{x\in[0,D]^{d}}\max_{1\leq i\leq d}|c_{i}(x)|.$
When $c(\cdot)$ is an invertible mapping, it belongs to the hidden convex optimization class. For $d=2$, $c_{1}(x)=x_{1}-1$, and $c_{2}(x)=x_{2}-2x_{1}^{2}+1$, we illustrate its contour plot in fig. 1.
Figure 1: The countour plots of the functions $F(x)=\frac{1}{4}(x_{1}-1)^{2}+\frac{1}{2}(2x_{1}^{2}-x_{2}-1)^{2}$ (top), and $F(x)=\max\left\{\frac{1}{4}|x_{1}-1|,\frac{1}{2}|2x_{1}^{2}-x_{2}-1|\right\}$ (bottom), $x=(x_{1},x_{2})^{\top}$ The left plots present the conour plots in the original space $\mathcal{X}$ and the right plots illustrate the reformulated space $\mathcal{U}$. The red star denotes the global minimum.
2.2 Minimizing posinomial functions [13, 29]
For power control in communication systems and optimal doping profile problems [13, 29], one often needs to minimize posinomial functions $F(\cdot):\mathbb{R}_{+}^{d}\rightarrow\mathbb{R}$ of the following form
$F(x)=\sum_{k=1}^{K}b_{k}x_{1}^{a_{1k}}\cdots x_{d}^{a_{dk}},$
where $b_{k}>0$ and $a_{ik}\in\mathbb{R}$ for all $k=1,\cdots,K$, $i=1,\cdots,d$. The function $F(\cdot)$ is non-convex but admits a convex reformulation via a variable change $u=c(x):=[\log(x_{1}),\ldots,\log(x_{d})]^{\top}$. The convex reformulation is of the form
$H(u)=F(c^{-1}(u))=\sum_{k=1}^{K}b_{k}e^{a_{1k}u_{1}}\cdots e^{a_{dk}u_{d}},$
where $H(\cdot)$ is convex. One can easily see that the above problem is hidden convex on a convex compact constraint $\mathcal{X}$.
2.3 System level synthesis in optimal control [2]
Consider a linear time-varying system
$x(t+1)=A_{t}\,x(t)+B_{t}\,u(t)+w(t),\qquad t=0,\ldots,T,$
where $x(t)\in\mathbb{R}^{n}$ is a state, $u(t)\in\mathbb{R}^{p}$ is a control input, and $w(t)\in\mathbb{R}^{n}$ is an exogenous disturbance process, and $x(0),w(t)\sim\mathcal{N}(0,\Sigma)$ are independent for $t=0,\ldots,T$. Matrices $A_{t}\in\mathbb{R}^{n\times n}$ and $B_{t}\in\mathbb{R}^{n\times p}$ determine the system dynamics. Define $\mathbf{x}=(x(0),\ldots,x(T))^{\top}$, $\mathbf{u}=(u(0),\ldots,u(T))^{\top}$, $\mathbf{w}=(x(0),w(0),\ldots,w(T-1))^{\top}$, and consider a time-varying controller of the form $u(t)=\sum_{i=0}^{t}K(t,t-i)x(i)$, which depends on a control matrix
$\mathbf{K}=\left[\begin{array}[]{cccc}K(0,0)&&&\\ K(1,1)&K(1,0)&&\\ \vdots&\ddots&\ddots&\\ K(T,T)&\cdots&K(T,1)&K(T,0)\end{array}\right].$
The goal of the system level synthesis is to find a control policy to minimize some loss functions, e.g., quadratic in $\mathbf{x}$ and $\mathbf{u}$: $F(\mathbf{K}):={\mathbb{E}}\left[\mathbf{x}^{\top}\mathcal{Q}\mathbf{x}+\mathbf{u}^{\top}\mathcal{R}\mathbf{u}\right],$ where $\mathcal{Q}=\text{diag}(Q_{0},\ldots,Q(T))\succeq 0$ and $\mathcal{R}=\text{diag}(R_{0},\ldots,R(T))\succeq 0$.
Despite the fact that $F(\cdot)$ is convex in both $\mathbf{x}$ and $\mathbf{u}$, it is non-convex in the decision variable $\mathbf{K}$. Nevertheless, it admits a convex reformulation [2] of the form
[…]
where $\Phi_{\mathbf{x}},\Phi_{\mathbf{u}}\in\mathbb{R}^{(T+1)\times(T+1)}$ are the new variables, $H(\cdot)$ is a strongly-convex function of $\Phi:=(\Phi_{\mathbf{x}},\Phi_{\mathbf{u}})$. $\mathbf{M}\in\mathbb{R}^{(T+1)\times(T+1)}$ is a deterministic matrix, which depends on matrices $A_{t}$, $B_{t}$, $t=0,\ldots,T-1$, and $I$ is the identity matrix. Moreover, there exists a bijection between variables $\mathbf{K}$ and $\Phi$ subject to the constraints of the reformulated problem. The (inverse of the) map $c(\cdot)$ is given by $\mathbf{K}=c^{-1}(\Phi):=\Phi_{\mathbf{u}}\Phi_{\mathbf{x}}^{-1}$ [2]. Therefore, one can easily verify that the optimization problem over $\mathbf{K}$ is hidden convex. A number of other problems in optimal control also admit suitable convex reformulations. We refer readers to [12, 80] for more examples.
2.4 Convex reinforcement learning [92]
Convex reinforcement learning (RL) problem generalizes the classical RL setting. It bases on a discounted Markov Decision Process $\mathbb{M}(\mathcal{S},\mathcal{A},\mathcal{P},H,\rho,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the (finite) state and action spaces respectively, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the state transition probability kernel (where $\Delta(\mathcal{S})$ denotes the distribution over $\mathcal{S}$), $\rho$ is the initial state distribution and $\gamma\in(0,1)$ is the discount factor. A stationary policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps each state $s\in\mathcal{S}$ to a distribution $\pi(\cdot|s)$ over the action space $\mathcal{A}$. The set of all (stationary) policies is denoted by $\Pi$ At each time step $h\in\mathbb{N}$ in a state $s_{h}\in\mathcal{S}$, the RL agent chooses an action $a_{h}\in\mathcal{A}$ with probability $\pi(a_{h}|s_{h})$ and the environment transitions to a state $s_{h+1}$ with probability $\mathcal{P}(s_{h+1}|s_{h},a_{h})\,.$ We denote by $\mathbb{P}_{\rho,\pi}$ the probability distribution of the Markov chain $(s_{h},a_{h})_{h\in\mathbb{N}}$ induced by the policy $\pi$ with an initial state distribution $\rho$. For any policy $\pi\in\Pi$, we define the state-action occupancy measure
$\lambda^{\pi}(s,a):=\sum_{h=0}^{+\infty}\gamma^{h}\mathbb{P}_{\rho,\pi}(s_{h}=s,a_{h}=a)\,\qquad\text{for all}~{}a\in\mathcal{A},\,s\in\mathcal{S}\,.$ (7)
The set of such state-action occupancy measures is denoted by $\mathcal{U}:=\{\lambda^{\pi}:\pi\in\Pi\}\,.$
Different from the classical RL, convex RL considers a general (convex) utility function $H:\mathcal{U}\to\mathbb{R}$ that maps the state-action occupancy measure to a cost and aims to find a policy that minimizes the cost
$\min_{\pi\in\Pi}F(\pi):=H(\lambda^{\pi}).$ (8)
Notice that $F(\cdot)$ is not convex in $\pi$ in general. However, for several commonly used utility functions, $H(\cdot)$ exhibits convexity in the occupancy measure $\lambda^{\pi}$. For standard RL, $H(\lambda^{\pi})=r^{\top}\lambda^{\pi}$ is linear in $\lambda^{\pi}$, where $r$ is the reward vector. For the pure exploration setting, focused on fully exploring the transitions in the environment, $H(\lambda^{\pi})$ represents the negative entropy of $\lambda^{\pi}$, which is also convex  [92]. For the imitation learning where the objective is to imitate the expert’s behavior given their sampled trajectories, $H(\lambda^{\pi})$ denotes the KL-divergence between $\lambda^{\pi}$ and the state-action occupancy measure learned from the expert’s sampled trajectories, which is also convex [92]. Thus, the convex RL problem belongs to the hidden convex class with $\mathcal{X}=\Pi$ and $c(x)=\lambda^{\pi}$ (with $x=\pi$). Under mild assumptions on the initial distribution $\rho$, the constant $\mu_{c}>0$ can be estimated, see e.g., [92]. Note that in convex RL, we can control $\lambda^{\pi}$ only implicitly by changing the policy $\pi$. The exact computation of the transformation map and its inverse requires the knowledge of the state transition probability kernel and can be computationally expensive.
2.5 Revenue management and inventory control [15, 16]
Consider a booking limit control in a passenger network revenue management problem. The goal is to maximize the revenue by finding an optimal booking limit threshold for each demand class, e.g., flying from New York to Seattle with economy class. Such a problem forms a two-stage stochastic programming such that
$\min_{x\in[0,D]^{d}}$ $\quad F(x):=\mathbb{E}_{\xi}[r^{\top}(x\wedge\xi)-\mathbb{E}_{\eta}\Gamma(x\wedge\xi,\eta)]$ (9)
$\mathrm{where}$ $\quad\Gamma(x\wedge\xi,\eta)=\min_{0\leq w\leq x\wedge\xi}\{l^{\top}(x\wedge\xi-w)\mid Aw\leq\eta\},$
where $d$ denotes the number of demand classes in the airline networks, $x\in\mathbb{R}^{d}$ is the booking limit control threshold for each demand class, $\xi$ is the random demand vector (of the same dimension as $x$) during the reservation stage, $x\wedge\xi$ denotes the number of reservations accepted, and $r^{\top}(x\wedge\xi)$ denotes revenue collected during the reservation stage with $r\in\mathbb{R}^{d}$ being the price vector. In the service stage, $\Gamma(x\wedge\xi,\eta)$ denotes the penalty on the airline companies when there are $x\wedge\xi$ number of reservations with plane seats capacity $\eta$ that is random, $w$ is the actual number of passengers that can get on the plane, $l$ is the penalty vector for declining passengers with reservation to get on the plane. Notice that $F$ is non-convex in $x$ due to the truncation between $x$ and $\xi$. However, when $\xi$ admits component-wise independent coordinates, this problem admits a convex reformulation via a variable change [15], i.e., $u=c(x)=\mathbb{E}_{\xi}[x\wedge\xi]$. Note that comparing to previous applications, the transformation function involves unknown distribution and thus is not explicitly known.
For more examples of hidden convex problems in operations research, we refer readers to [36] about supply chain management and [16, 61] about revenue management.
3 Properties of Hidden Convex Optimization
In this section, we provide key properties of hidden convex problems and discuss its connections with gradient dominated function classes.
3.1 Globally optimal solution
The following proposition suggests that every stationary point of a hidden convex function is a global minima.
Proposition 1.
Let $F(\cdot)$ be hidden convex on $\mathcal{X}$ and $\bar{x}\in\mathcal{X}$ be a stationary point. If the map $c(\cdot)$ is differentiable at $\bar{x}$ , then $\bar{x}$ is a global minimum, i.e., $F(\bar{x})\leq F(x)$ for any $x\in\mathcal{X}$.
Proof.
By the definition of a stationary point and the chain rule [74] (Theorem 10.49), we can write
$0\in\partial_{x}(F+\delta_{\mathcal{X}})(\bar{x})=\nabla c(\bar{x})\left(\partial_{u}H(\bar{u})+\partial_{u}\delta_{\mathcal{U}}(\bar{u})\right),$ (10)
where $\bar{u}=c(\bar{x})$. As the map $c(\cdot)$ is invertible, then $\nabla c(\bar{x})y=0$ for some $y\in\mathbb{R}^{d}$ implies $y=0$. Thus, we have $0\in\partial_{u}H(\bar{u})+\partial_{u}\delta_{\mathcal{U}}(\bar{u})$. Since function $H(\cdot)$ is convex, by the sufficient optimality condition, $\bar{u}$ is a globally optimal solution, i.e., $H(\bar{u})\leq H(u)$ for any $u\in\mathcal{U}$. As a result, we have $F(\bar{x})=H(\bar{u})\leq H(u)=F(x)$ for any $x\in\mathcal{X}$. ∎
Note that a similar result appeared in [92] under additional smoothness assumptions on $H(\cdot)$ and $c(\cdot)$. The above proof is much simpler and does not require smoothness.
3.2 Connections with gradient dominated functions
It is natural to ask what is the connection between hidden convex problems and previously studied gradient dominated function classes that also ensure the global convergence of gradient-based algorithms in the smooth setting [47]. Unfortunately, the exact characterization is difficult to establish in the constrained setting. With the following proposition, we show that a problem satisfies the global KŁ condition if it is hidden strongly convex ($\mu_{H}>0$).
Proposition 2.
Let $F(\cdot)$ be differentiable and $(\mu_{c},\mu_{H})$-hidden strongly convex, the map $c(\cdot)$ be differentiable on $\mathcal{X}$. Then the optimization problem satisfies the global KŁ condition, i.e.,
$\min_{h_{x}\in\partial\delta_{\mathcal{X}}(x)}\left\|\nabla F(x)+h_{x}\right\|^{2}\geq 2\mu_{H}\mu_{c}^{2}\left(F(x)-F^{*}\right)\qquad\text{for all }x\in\mathcal{X}.$ (11)
Proof.
Since $H(\cdot)$ is differentiable and strongly convex on a convex set $\mathcal{U}$, then it satisfies the following Proximal-PŁ condition [47] (Appendix F)
$\mathcal{D_{\delta_{\mathcal{U}}}}(u,\mu_{H})\geq 2\mu_{H}(H(u)-H(u^{*}))\qquad\text{for all }u\in\mathcal{U},$
where $\mathcal{D_{\delta_{\mathcal{U}}}}(u,\mu_{H}):=-2\mu_{H}\min_{v\in\mathcal{U}}\left\{\langle\nabla H(u),v-u\rangle+\frac{\mu_{H}}{2}\left\|u-v\right\|^{2}\right\}$. The inequality above implies global KŁ condition, see Appendix G in [47],
$\min_{h_{u}\in\partial\delta_{\mathcal{U}}(u)}\left\|\nabla H(u)+h_{u}\right\|^{2}\geq 2\mu_{H}\left(H(u)-H(u^{*})\right).$
Combined with the chain rule, we have for any $u=c(x)$
$\min_{h_{x}\in\partial\delta_{\mathcal{X}}(x)}\left\|\nabla F(x)+h_{x}\right\|^{2}$ $=$ $\min_{h_{u}\in\partial\delta_{\mathcal{U}}(u)}\left\|\nabla c(x)(\nabla H(u)+h_{u})\right\|^{2}\geq 2\mu_{H}\mu_{c}^{2}\left(F(x)-F^{*}\right),$
where $F^{*}=F(x^{*})$. This concludes the proof. ∎
Note that even in this restrictive case when $\mu_{H}>0$, it remains elusive how condition (11) can be used to establish global convergence of Projected SGD. To our knowledge, under the condition (11), no global analysis of stochastic gradient methods exists in the literature even for smooth $F(\cdot)$. In the more interesting case when $H(\cdot)$ is merely convex, the above global KŁ condition becomes vacuous as $\mu_{H}=0$. Despite of this, Proposition 1 still holds even when $\mu_{H}=0$.
3.3 Key inequalities for analysis of gradient methods
The following observations are key tools for deriving global convergence under hidden convexity.
Proposition 3.
Let $F(\cdot)$ be hidden convex with $\mu_{c}>0,\mu_{H}\geq 0$. For any $\alpha\in[0,1]$, $x^{*}\in\mathcal{X}^{*}$ and $x\in\mathcal{X}$ , define $x_{\alpha}:=c^{-1}\left((1-\alpha)c(x)+\alpha c(x^{*})\right)$. Then
$F(x_{\alpha})\leq(1-\alpha)F(x)+\alpha F(x^{*})-\frac{(1-\alpha)\alpha\mu_{H}}{2}\left\|c(x)-c(x^{*})\right\|^{2},$ (12)
and
$\left\|x_{\alpha}-x\right\|\leq\frac{\alpha}{\mu_{c}}\left\|c(x)-c(x^{*})\right\|.$ (13)
Proof.
By (strong) convexity of $H(\cdot)$ and convexity of $\mathcal{U}$, we have
$F(x_{\alpha})$ $=$ $F(c^{-1}\left((1-\alpha)c(x)+\alpha c(x^{*})\right))$
$=$ $H((1-\alpha)c(x)+\alpha c(x^{*}))$
$\leq$ $(1-\alpha)H(c(x))+\alpha H(c(x^{*}))-\frac{(1-\alpha)\alpha\mu_{H}}{2}\left\|c(x)-c(x^{*})\right\|^{2}$
$=$ $(1-\alpha)F(x)+\alpha F(x^{*})-\frac{(1-\alpha)\alpha\mu_{H}}{2}\left\|c(x)-c(x^{*})\right\|^{2}.$
where the inequality uses the fact that $\mathcal{U}$ is a convex set and that $(1-\alpha)c(x)+\alpha c(x^{*})\in\mathcal{U}$ for any $x\in\mathcal{X}$. By definition of $x_{\alpha}$ and (6), we derive
$\left\|x_{\alpha}-x\right\|=\left\|c^{-1}\left((1-\alpha)c(x)+\alpha c(x^{*})\right)-c^{-1}(c(x))\right\|\leq\frac{1}{\mu_{c}}\left\|\alpha(c(x)-c(x^{*}))\right\|.$
4 Stochastic Subgradient Method
In this section, we show how Proposition 3 can be used to analyze convergence of the projected stochastic subgradient method (SM) as described in (3) in the non-smooth setting.
We first make the following assumptions.
  1. A.1
    $F(\cdot)$ is $\ell$-weakly convex on a closed, convex set $\mathcal{X}$.
  2. A.2
    We have access to a stochastic sub-gradient oracle of $F(\cdot)$ at any $x\in\mathcal{X}$, which outputs a random vector $g(x,\xi)$ such that ${\mathbb{E}}\left[g(x,\xi)\right]\in\partial F(x),$ where $\partial F(x)$ is the sub-differential set of $F(\cdot)$ at $x$. Moreover, there exists $G_{F}>0$
    ${\mathbb{E}}\left[\left\|g(x,\xi)\right\|^{2}\right]\leq G_{F}^{2}\qquad\text{for any }x\in\mathcal{X}.$
The above assumptions are standard and appear frequently in non-smooth optimization [24, 95, 23, 22]. Weak convexity is known to be a much weaker condition than smoothness [22]. Notably, in the context of our hidden convexity (C.1. and C.2.), weak convexity is not restrictive and comes for free from the Lipschitz continuity of $H(\cdot)$ and the smoothness of the transformation function $c(\cdot)$. Specifically, if $H:\mathcal{U}\rightarrow\mathbb{R}$ is convex and $G_{H}$-Lipschitz continuous on $\mathcal{U}$ and $c:\mathcal{X}\rightarrow\mathcal{U}$ is $L_{c}$-smooth, then it can be shown that the composition $F(x)=H(c(x))$ is $\ell$-weakly convex with $\ell:=G_{H}L_{c}$; see e.g. Proposition 2.2(c) in [95]. In the absence of smoothness, the second assumption on bounded second moment of the (stochastic) sub-gradients is typical even in convex case. Later in Section 5, we show this assumption can be further relaxed to bounded variance in the smooth setting.
Let $x^{*}\in\mathcal{X}^{*}$, and $\Phi:=F+\delta_{\mathcal{X}}$. We define the Lyapunov function
$\Lambda_{t}:={\mathbb{E}}\left[\Phi_{1/\rho}(x^{t})-F(x^{*})\right],$
where $\Phi_{1/\rho}$ is the Moreau envelope of $\Phi$. Notice that $\Lambda_{t}\geq 0$ for any $t\geq 0$ and $\Lambda_{t}=0$ if and only if $x^{t}\in\mathcal{X}^{*}$.
Before stating the main result, we recall the following useful lemma from [22] that controls the distance between one step of the SM, $x^{t+1}$, and one step of the proximal point method, $\hat{x}^{t}$. We include its proof in Appendix A for completeness.
Lemma 1 (Lemma 3.3 in [22]).
Let A.1, A.2 hold, and $\rho>\ell$, $\eta\leq 1/\rho$ Then for all $t\geq 0$, we have $\mathbb{E}\left[\left\|x^{t+1}-\hat{x}^{t}\right\|^{2}\mid x^{t}\right]\leq(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+2G_{F}^{2}\eta^{2}.$
The next theorem is the essential step for establishing the global convergence of SM in theorems 2 and 3.
Theorem 1.
Let C.1, C.2, A.1 and A.2 hold with $\mu_{H}\geq 0$. Set $\rho=2\ell$, $\eta\leq\frac{1}{2\ell}$. Define $\hat{x}^{t}:=\operatorname{prox}_{\Phi/\rho}(x^{t})$. Then for any $0<\alpha\leq\eta\ell$ and $t\geq 0$
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+\left(\frac{3\alpha^{2}}{2\mu_{c}^{2}\eta}-\frac{(1-\alpha)\alpha\mu_{H}}{2}\right){\mathbb{E}}\left[\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}\right]+4\ell\eta^{2}G_{F}^{2}.$
Proof.
By the definition of $\hat{x}^{t+1}$, we have for any $z\in\mathcal{X}$
$\mathbb{E}\left[\Phi_{1/\rho}\left(x^{t+1}\right)\right]$ $=\mathbb{E}\left[\Phi\left(\hat{x}^{t+1}\right)+\frac{\rho}{2}\left\|\hat{x}^{t+1}-x^{t+1}\right\|^{2}\right]$
$\overset{(i)}{\leq}\mathbb{E}\left[\Phi\left(z\right)+\frac{\rho}{2}\left\|z-x^{t+1}\right\|^{2}\right]$
$\overset{(ii)}{\leq}\mathbb{E}\left[\Phi\left(z\right)+\left(1+s\right)\frac{\rho}{2}\left\|\hat{x}^{t}-x^{t+1}\right\|^{2}+\left(1+\frac{1}{s}\right)\frac{\rho}{2}\left\|\hat{x}^{t}-z\right\|^{2}\right]$
$\overset{(iii)}{\leq}\mathbb{E}\left[\Phi\left(z\right)+\left(1+s\right)(1-\eta\rho)\frac{\rho}{2}\left\|\hat{x}^{t}-x^{t}\right\|^{2}\right]$
$\qquad+\left(1+\frac{1}{s}\right)\frac{\rho}{2}{\mathbb{E}}\left[\left\|\hat{x}^{t}-z\right\|^{2}\right]+\left(1+s\right)\rho\eta^{2}G_{F}^{2},$
where in $(i)$ we use the optimality of $\hat{x}^{t+1}$, $(ii)$ follows from Young’s inequality for any $s>0$, and in $(iii)$ we apply the result of lemma 1. We now select $s=\eta\rho/2$, which guarantees $\left(1+s\right)(1-\eta\rho)\leq 1-\eta\rho/2$, $1+s\leq 2$, and $1+1/s\leq 3/(\eta\rho)$. Thus
$\mathbb{E}\left[\Phi_{1/\rho}\left(x^{t+1}\right)\right]$$\leq$ $\mathbb{E}\left[F\left(z\right)+\left(1-\frac{\eta\rho}{2}\right)\frac{\rho}{2}\left\|\hat{x}^{t}-x^{t}\right\|^{2}\right]+\frac{3}{2\eta}{\mathbb{E}}\left[\left\|\hat{x}^{t}-z\right\|^{2}\right]+2\rho\eta^{2}G_{F}^{2}.$
We are now ready to utilize the properties of hidden convex functions to bound $F(z)$ and $\left\|\hat{x}^{t}-z\right\|^{2}$ for some specific choice of $z\in\mathcal{X}$. By Proposition 3, we have for $z=\hat{x}_{\alpha}^{t}:=c^{-1}((1-\alpha)c(\hat{x}^{t})+\alpha c(x^{*}))$
$F(z)\leq(1-\alpha)F(\hat{x}^{t})+\alpha F(x^{*})-\frac{(1-\alpha)\alpha\mu_{H}}{2}\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2},$
$\left\|z-\hat{x}^{t}\right\|^{2}\leq\frac{\alpha^{2}}{\mu_{c}^{2}}\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}.$
Combining three inequalities above, we have
${\mathbb{E}}\left[\Phi_{1/\rho}(x^{t+1})\right]$$\leq$ $(1-\alpha){\mathbb{E}}\left[F(\hat{x}^{t})\right]+\alpha F(x^{*})+\left(1-\frac{\eta\rho}{2}\right)\frac{\rho}{2}{\mathbb{E}}\left[\left\|\hat{x}^{t}-x^{t}\right\|^{2}\right]+2\rho\eta^{2}G_{F}^{2}$
$\qquad+\left(\frac{3\alpha^{2}}{2\mu_{c}^{2}\eta}-\frac{(1-\alpha)\alpha\mu_{H}}{2}\right){\mathbb{E}}\left[\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}\right]$
$\leq$ $(1-\alpha){\mathbb{E}}\left[\Phi_{1/\rho}(x^{t})\right]+\alpha F(x^{*})+2\rho\eta^{2}G_{F}^{2}$
$\qquad+\left(\frac{3\alpha^{2}}{2\mu_{c}^{2}\eta}-\frac{(1-\alpha)\alpha\mu_{H}}{2}\right){\mathbb{E}}\left[\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}\right],$
where the last inequality holds since $1-\frac{\eta\rho}{2}\leq 1-\alpha$ (by the choice $\alpha\leq\eta\ell$, $\rho=2\ell$) and recognizing $\Phi_{1/\rho}(x^{t})$. Subtracting $F(x^{*})$ from both sides, we conclude the proof. ∎
4.1 Hidden Convex Setting
We first demonstrate the convergence rate of SM in the hidden convex setting.
Theorem 2.
Let C.1, C.2, A.1, A.2 hold with $\mu_{H}=0$, and the set $\mathcal{U}$ be bounded by a diameter $D_{\mathcal{U}}$. Fix $\varepsilon>0$, and set the step-size in (3) as $\eta=\frac{1}{2\ell}\cdot\min\left\{1,\frac{\mu_{c}^{2}\varepsilon^{2}}{24D_{\mathcal{U}}^{2}G_{F}^{2}}\right\}.$ Then for $\rho=2l$, we have $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{\ell D_{\mathcal{U}}^{2}}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{\ell D_{\mathcal{U}}^{4}G_{F}^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right)$ iterations.
Proof.
Setting $\mu_{H}=0$ in Theorem 1 and leveraging compactness of $\mathcal{U}$, we have
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+\frac{3D_{\mathcal{U}}^{2}\alpha^{2}}{2\mu_{c}^{2}\eta}+4\ell\eta^{2}G_{F}^{2}.$
Unrolling the recursion for $t=0$ to $t=T-1$, we get
$\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{3D_{\mathcal{U}}^{2}\alpha}{2\mu_{c}^{2}\eta}+\frac{4\ell\eta^{2}G_{F}^{2}}{\alpha}\leq\varepsilon,$
where the last step holds by setting $\alpha=\min\left\{\eta\ell,\frac{2\varepsilon\mu_{c}^{2}\eta}{3D_{\mathcal{U}}^{2}},\frac{\sqrt{16\ell}\mu_{c}G_{F}\eta^{\nicefrac{{3}}{{2}}}}{\sqrt{3}D_{\mathcal{U}}}\right\}$ after $T=\frac{1}{\alpha}\log\left(\frac{3\Lambda_{0}}{\varepsilon}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\varepsilon}+\frac{\ell D_{\mathcal{U}}^{4}G_{F}^{2}}{\mu_{c}^{4}\varepsilon^{3}}\right).$
We remark that in the absence of smoothness of $F(\cdot)$, the guarantee on $\Lambda_{t}$ might not necessarily translate to the function value gap $F(x^{t})-F(x^{*})$. However, with the following corollary we show that the output of SM, $x^{T}$, is in fact close to an $\varepsilon$-approximate global solution $\hat{x}^{T}=\operatorname{prox}_{\Phi/\rho}(x^{T})$.
Corollary 1.
Under the setting of Theorem 2, SM finds a point $x^{T}\in\mathcal{X}$, which is close to $\hat{x}^{T}$, an $\varepsilon$-global solution of (1). More specifically, it holds that ${\mathbb{E}}\left[\left\|\hat{x}^{T}-x^{T}\right\|^{2}\right]\leq\varepsilon/\ell$ and ${\mathbb{E}}\left[F(\hat{x}^{T})-F(x^{*})\right]\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}(\varepsilon^{-3})$.
Proof.
The result follows directly from the definition of $\Lambda_{T}$ and Theorem 2. ∎
4.2 Hidden Strongly Convex Setting
The following theorem presents a stronger result in the case when $F(\cdot)$ is additionally hidden strongly convex.
Theorem 3.
Let C.1, C.2, A.1, A.2 hold with $\mu_{H}>0$. Then for $\rho=2\ell$ and any $\eta\leq\frac{1}{2\ell}$, $\alpha\leq\min\left\{\eta\ell,\frac{\eta\mu_{c}^{2}\mu_{H}}{2}\right\}$, we have for all $T\geq 0$ $\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{4\ell\eta^{2}G_{F}^{2}}{\alpha}.$ Fix $\varepsilon>0$, and set the step-size in (15) as $\eta=\min\left\{\frac{1}{2\ell},\frac{\mu_{c}^{2}\mu_{H}\varepsilon}{10\ell G_{F}^{2}}\right\}.$ Then $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{\ell}{\mu_{c}^{2}\mu_{H}}+\frac{\ell G_{F}^{2}}{\mu_{c}^{4}\mu_{H}^{2}}\frac{1}{\varepsilon}\right)$ iterations.
Proof.
We invoke Theorem 1 with $\mu_{H}>0$. The choice of $\alpha$ guarantees the coefficient in front of ${\mathbb{E}}\left[\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}\right]$ is non-positive and
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+4\ell\eta^{2}G_{F}^{2}.$
It remains to conclude the proof by unrolling the recursion and setting the step-size accordingly. ∎
In the presence of hidden strong convexity, since the optimal $x^{*}\in\mathcal{X}^{*}$ is unique, we can establish a strong convergence of the sequence $\{x^{t}\}_{t\geq 0}$ to $x^{*}$.
Corollary 2.
Let the assumptions of Theorem 3 hold and $x^{T}$ be the output of the method (3) after $T$ iterations (given by Theorem 3). Then ${\mathbb{E}}\left[\left\|x^{T}-x^{*}\right\|^{2}\right]\leq\left(\frac{4}{\mu_{H}\mu_{c}^{2}}+\frac{2}{\ell}\right)\varepsilon$
Proof.
Since $H(\cdot)$ is $\mu_{H}$-strongly convex on $\mathcal{X}$, we have
$F(\hat{x}^{T})-F(x^{*})=H(c(\hat{x}^{T}))-H(c(x^{*}))\geq\frac{\mu_{H}}{2}\left\|c(\hat{x}^{T})-c(x^{*})\right\|^{2}\geq\frac{\mu_{H}\mu_{c}^{2}}{2}\left\|\hat{x}^{T}-x^{*}\right\|^{2},$
where the first inequality follows by the first-order characterization of strong convexity and the optimality condition, and the last inequality holds by C.2. Recall that $\Lambda_{T}={\mathbb{E}}\left[F(\hat{x}^{T})-F(x^{*})+\frac{\rho}{2}\left\|\hat{x}^{T}-x^{T}\right\|^{2}\right]$ with $\rho=2\ell$. Then
${\mathbb{E}}\left[\left\|x^{T}-x^{*}\right\|^{2}\right]$$\leq$ $2{\mathbb{E}}\left[\left\|\hat{x}^{T}-x^{*}\right\|^{2}\right]+2{\mathbb{E}}\left[\left\|\hat{x}^{T}-x^{T}\right\|^{2}\right]$
$\leq$ $\frac{4}{\mu_{H}\mu_{c}^{2}}{\mathbb{E}}\left[F(\hat{x}^{T})-F(x^{*})\right]+2\,{\mathbb{E}}\left[\left\|\hat{x}^{T}-x^{T}\right\|^{2}\right]\leq\left(\frac{4}{\mu_{H}\mu_{c}^{2}}+\frac{2}{\ell}\right)\varepsilon.$
where the second inequality holds by (4.2) and the last step follows by Theorem 3. ∎
The above results highlight a notable observation: although $F(\cdot)$ is non-smooth and non-convex, simple SM converges to a globally optimal solution. This stands in contrast to recent results in general non-smooth non-convex optimization, where more sophisticated (randomized) algorithms are needed to obtain a meaningful solution (e.g., $(\delta,\epsilon)$-Goldstein stationary point) [46]. Moreover, it is worth emphasizing the distinctions in the sample complexity results compared to classical findings in convex settings: Theorems 2 and 3 implies the sample complexities of $\widetilde{\mathcal{O}}(\varepsilon^{-3})$ and $\widetilde{\mathcal{O}}(\varepsilon^{-1})$ respectively to reach ${\mathbb{E}}\left[\Phi_{1/\rho}(x)-F(x^{*})\right]\leq\varepsilon$ for hidden convex and hidden strongly convex problems, whereas in convex and strongly convex settings, the sample complexities are ${\mathcal{O}}(\varepsilon^{-2})$ and ${\mathcal{O}}(\varepsilon^{-1})$ respectively to reach ${\mathbb{E}}\left[F(x)-F^{*}\right]\leq\varepsilon$ [64].
5 Projected SGD
In this section, we consider the smooth setting when $F$ is continuously differentiable. In this case, SM reduces to Projected SGD (P-SGD):
$x^{t+1}=\Pi_{\mathcal{X}}(x^{t}-\eta\nabla f(x^{t},\xi^{t})).$ (15)
In particular, we assume that
  1. A.1’
    The function $F:\mathcal{X}\rightarrow\mathbb{R}$ is differentiable on a closed, convex set $\mathcal{X}$ and its gradient $\nabla F(x)$ is $L$-Lipschitz continuous.
  2. A.2’
    We have access to an unbiased stochastic gradient oracle with bounded variance $\sigma>0$, i.e. for any $x\in\mathcal{X}$: ${\mathbb{E}}\left[\nabla f(x,\xi)\right]=\nabla F(x)$, and
    ${\mathbb{E}}\left[\left\|\nabla f(x,\xi)-\nabla F(x)\right\|^{2}\right]\leq\sigma^{2},$
    where expectations are with respect to the random variable $\xi\sim\mathcal{D}$.
Note that Assumption A.2’ on bounded variance is considerably weaker than Assumption A.2 on the bounded second moment of stochastic gradient in previous section. Replacing lemma 1 with lemma 4 in the proof of theorem 1 of the previous section, we are able to derive the following results under Assumptions A.1’ and A.2’.
Theorem 4.
Let C.1, C.2, A.1’, A.2’ hold with $\mu_{H}\geq 0$. Set $\rho=4L$, $\eta\leq\frac{2}{9L}$. Define $\hat{x}^{t}:=\operatorname{prox}_{\Phi/\rho}(x^{t})$. Then for any $0<\alpha\leq 2\eta L$ and $t\geq 0$
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+\rho\eta^{2}\sigma^{2}+\left(\frac{3\alpha^{2}}{2\mu_{c}^{2}\eta}-\frac{(1-\alpha)\alpha\mu_{H}}{2}\right){\mathbb{E}}\left[\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}\right].$
Using the above result, we provide a refined analysis of Projected SGD in the differentiable setting with smoothness and bounded variance.
5.1 Hidden Convex Setting
We start with the hidden convex case.
Theorem 5.
Let C.1, C.2, A.1’, A.2’ hold with $\mu_{H}=0$, and the set $\mathcal{U}$ be bounded by a diameter $D_{\mathcal{U}}$. Then for $\rho=4L$ and any $\eta\leq\frac{2}{9L}$, $\alpha\leq 2\eta L$, we have for all $T\geq 0$ $\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{3D_{\mathcal{U}}^{2}\alpha}{2\mu_{c}^{2}\eta}+\frac{4L\eta^{2}\sigma^{2}}{\alpha}.$ Fix $\varepsilon>0$, and set the step-size in (15) as $\eta=\frac{2}{9L}\cdot\min\left\{1,\frac{\mu_{c}^{2}\varepsilon^{2}}{12D_{\mathcal{U}}^{2}\sigma^{2}}\right\}.$ Then $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{LD_{\mathcal{U}}^{2}}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{LD_{\mathcal{U}}^{4}\sigma^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right)$ iterations.
Proof.
We set $\alpha=\min\left\{2\eta L,\frac{2\varepsilon\mu_{c}^{2}\eta}{3D_{\mathcal{U}}^{2}},\frac{\sqrt{8L}\mu_{c}\sigma\eta^{\nicefrac{{3}}{{2}}}}{\sqrt{3}D_{\mathcal{U}}}\right\}$, then, given Theorem 4, the reminder of the proof is similar to the proof of Theorem 2 in the previous section. ∎
Similar to Corollary 1, we can show that $x^{T}$ is close to an $\varepsilon$-global optimal solution. But in the case of smooth $F(\cdot)$, we can also derive a stronger result after applying one (post-processing) step of Projected SGD with mini-batch. We defer this result to Corollary 4 in Appendix B.
Theorem 5 implies that in deterministic case when $\sigma^{2}=0$, the iteration complexity of the gradient method is $\widetilde{\mathcal{O}}(\varepsilon^{-1})$, which coincides with the iteration complexity of Projected GD in the smooth convex setting in terms of $\varepsilon$ (up to a logarithmic factor). However, in the stochastic setting, $\widetilde{\mathcal{O}}(\varepsilon^{-3})$ sample complexity is worse than the well known ${\mathcal{O}}(\varepsilon^{-2})$ sample complexity in the convex case [51]. On the other hand, for general smooth nonconvex optimization, Projected SGD is only known to converge to a first-order stationary point (FOSP), i.e., find $x\in\mathcal{X}$ with ${\mathbb{E}}\left[\left\|\nabla F(x)\right\|\right]\leq\epsilon$, with the sample complexity ${\mathcal{O}}(\epsilon^{-4})$ [51, 22].
5.2 Hidden Strongly Convex Setting
Similar to the exposition in Section 4, we present an improved sample complexity result for hidden strongly convex problems.
Theorem 6.
Let C.1, C.2, A.1’, A.2’ hold with $\mu_{H}>0$. Then for $\rho=4L$ and any $\eta\leq\frac{2}{9L}$, $\alpha\leq\min\left\{2\eta L,\frac{\eta\mu_{c}^{2}\mu_{H}}{2}\right\}$, we have for all $T\geq 0$ $\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{4L\eta^{2}\sigma^{2}}{\alpha}.$ Fix $\varepsilon>0$, and set the step-size in (15) as $\eta=\min\left\{\frac{2}{9L},\frac{\mu_{c}^{2}\mu_{H}\varepsilon}{10L\sigma^{2}}\right\}.$ Then $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{L}{\mu_{c}^{2}\mu_{H}}+\frac{L\sigma^{2}}{\mu_{c}^{4}\mu_{H}^{2}}\frac{1}{\varepsilon}\right)$ iterations.
Proof.
The proof follows from Theorem 4 using the same steps as in the proof of Theorem 3. ∎
Similarly to corollary 2, we can translate convergence in $\Lambda_{T}$ to the last iterate convergence in terms of distance to the optimal solution.
Corollary 3.
Let the assumptions of Theorem 6 hold and $x^{T}$ be the output of the method (15) after $T$ iterations (given by Theorem 6). Then ${\mathbb{E}}\left[\left\|x^{T}-x^{*}\right\|^{2}\right]\leq\left(\frac{4}{\mu_{H}\mu_{c}^{2}}+\frac{1}{L}\right)\varepsilon$.
Theorem 6 and Corollary 3 imply that if $\mu_{H}>0$, Projected SGD converges linearly in deterministic setting (when $\sigma=0$) and achieves $\widetilde{\mathcal{O}}(\varepsilon^{-1})$ sample complexity in the stochastic setting. This means that compared to the special case of strongly convex optimization, the above rates have the same dependence on $\varepsilon$ (up to a logarithmic factor) [79, 51].
6 Projected SGD with Momentum
We observe that the previous section only guarantees convergence of $\Lambda_{t}$, however, this might not directly imply the convergence on the original function $F(\cdot)$ since $\Phi_{1/\rho}(x)\leq F(x)$ for any $x\in\mathcal{X}$. It is known that in convex optimization, momentum is often helpful to establish the last iterate convergence, see e.g., [55, 76]. Motivated by this, we consider Projected SGD with Polyak’s (heavy-ball) momentum [69] in the smooth setting. We show that with extra momentum step, we can establish last-iterate convergence to an $\varepsilon$-optimal solution. The Projected SGD with momentum admits the following updates:
$x^{t+1}=\Pi_{\mathcal{X}}(x^{t}-\eta\,g^{t}),\qquad g^{t+1}=(1-\beta)\,g^{t}+\beta\,\nabla f(x^{t+1},\xi^{t+1}).$ (16)
Our analysis in this section uses the same properties presented in Section 3.3, but the Lyapunov function used here is completely different from $\Lambda_{t}$ used in Sections 4 and 5. Let $x^{*}\in\mathcal{X}^{*}$, for any $x^{t}\in\mathcal{X}$, we define the Lyapunov function
$\Lambda_{t}^{HB}:={\mathbb{E}}\left[F(x^{t})-F(x^{*})+\frac{\eta}{\beta}\left\|g^{t}-\nabla F(x^{t})\right\|^{2}\right].$ (17)
The following lemma controls the error between the momentum gradient estimator $g^{t}$ and the true gradient $\nabla F(x^{t})$. Similar recursive error control was previously used in general non-convex optimization, e.g., in [21, 19, 34].
Lemma 2 (Lemma 2 in [34]).
Let $\beta\in(0,1]$ and $g^{t}$ be updated via (16). Then
${\mathbb{E}}\left[\left\|g^{t+1}-\nabla F(x^{t+1})\right\|^{2}\right]\leq(1-\beta){\mathbb{E}}\left[\left\|g^{t}-\nabla F(x^{t})\right\|^{2}\right]+\frac{3L^{2}}{\beta}{\mathbb{E}}\left[\left\|x^{t}-x^{t+1}\right\|^{2}\right]+\beta^{2}\sigma^{2}.$
The following result is the key to derive global convergence guarantee for Projected SGD with momentum under hidden convexity.
Theorem 7.
Suppose that C.1, C.2, A.1’, A.2’ hold with $\mu_{H}\geq 0$, and the step-size in (16) satisfies $\eta\leq 1/L$. For any $\alpha\in[0,1]$, it holds that
$F(x^{t+1})$$\leq$ $(1-\alpha)F(x^{t})+\alpha F(x^{*})+\left(\frac{\alpha^{2}}{\mu_{c}^{2}\eta}-\frac{(1-\alpha)\alpha\mu_{H}}{2}\right)\left\|c(x^{t})-c(x^{*})\right\|^{2},$ (18)
$\qquad-\left(\frac{1}{2\eta}-\frac{L}{2}\right)\left\|x^{t+1}-x^{t}\right\|^{2}+\frac{\eta}{2}\left\|g^{t}-\nabla F(x^{t})\right\|^{2}.$
Proof.
By the update rule of $x^{t+1}$ and following the standard descent inequality (cf. lemma 5), we have for any $z\in\mathcal{X}$ that
$\langle g^{t},x^{t+1}-z\rangle+\frac{1}{2\eta}\left\|x^{t+1}-x^{t}\right\|^{2}\leq\frac{1}{2\eta}\left\|z-x^{t}\right\|^{2}-\frac{1}{2\eta}\left\|z-x^{t+1}\right\|^{2}.$ (19)
By the smoothness of $F(\cdot)$, we derive
$F(x^{t+1})$$\leq$ $F(x^{t})+\langle\nabla F(x^{t}),x^{t+1}-x^{t}\rangle+\frac{L}{2}\left\|x^{t+1}-x^{t}\right\|^{2}$
$=$ $F(x^{t})+\langle g^{t},x^{t+1}-x^{t}\rangle+\frac{1}{2\eta}\left\|x^{t+1}-x^{t}\right\|^{2}$
$\qquad+\langle\nabla F(x^{t})-g^{t},x^{t+1}-x^{t}\rangle-\left(\frac{1}{2\eta}-\frac{L}{2}\right)\left\|x^{t+1}-x^{t}\right\|^{2}$
$\overset{(i)}{\leq}$ $F(x^{t})+\langle g^{t},z-x^{t}\rangle+\frac{1}{2\eta}\left\|z-x^{t}\right\|^{2}-\frac{1}{2\eta}\left\|z-x^{t+1}\right\|^{2}$
$\qquad+\langle\nabla F(x^{t})-g^{t},x^{t+1}-x^{t}\rangle-\left(\frac{1}{2\eta}-\frac{L}{2}\right)\left\|x^{t+1}-x^{t}\right\|^{2}$
[…]
$\qquad+\langle\nabla F(x^{t})-g^{t},x^{t+1}-z\rangle-\left(\frac{1}{2\eta}-\frac{L}{2}\right)\left\|x^{t+1}-x^{t}\right\|^{2}$
$\overset{(ii)}{\leq}$ $F(x^{t})+\langle\nabla F(x^{t}),z-x^{t}\rangle+\frac{1}{2\eta}\left\|z-x^{t}\right\|^{2}+\frac{\eta}{2}\left\|g^{t}-\nabla F(x^{t})\right\|^{2}$
$\qquad-\left(\frac{1}{2\eta}-\frac{L}{2}\right)\left\|x^{t+1}-x^{t}\right\|^{2}$
$\overset{(iii)}{\leq}$ $F(z)+\frac{L}{2}\left\|z-x^{t}\right\|^{2}+\frac{1}{2\eta}\left\|z-x^{t}\right\|^{2}+\frac{\eta}{2}\left\|g^{t}-\nabla F(x^{t})\right\|^{2}$
$\qquad-\left(\frac{1}{2\eta}-\frac{L}{2}\right)\left\|x^{t+1}-x^{t}\right\|^{2},$
where $(i)$ follows from (19), $(ii)$ holds by Young’s inequality, i.e., $\langle a,b\rangle\leq\frac{\eta}{2}\left\|a\right\|^{2}+\frac{1}{2\eta}\left\|b\right\|^{2}$ with $a=\nabla F(x^{t})-g^{t}$, $b=x^{t+1}-z$, $(iii)$ holds by the smoothness of $F(\cdot)$, i.e., $-\frac{L}{2}\left\|z-x^{t}\right\|^{2}\leq F(x^{t})-F(z)-\langle\nabla F(x^{t}),z-x^{t}\rangle$. We are now ready to utilize the properties of hidden convex functions to bound $F(z)$ and $\left\|z-x^{t}\right\|^{2}$ for some specific choice of $z\in\mathcal{X}$. We select $z:=x_{\alpha}^{t}=c^{-1}((1-{\alpha})c(x^{t})+{\alpha}c(x^{*}))\in\mathcal{X}$, for some $\alpha\in[0,1]$, and $x^{*}\in\mathcal{X}^{*}$. By Proposition 3, we have for $\mu_{H}\geq 0$,
$F(z)\leq(1-\alpha)F(x^{t})+\alpha F(x^{*})-\frac{(1-\alpha)\alpha\mu_{H}}{2}\left\|c(x^{t})-c(x^{*})\right\|^{2},$
and
$\left\|z-x^{t}\right\|^{2}\leq\frac{\alpha^{2}}{\mu_{c}^{2}}\left\|c(x^{t})-c(x^{*})\right\|^{2}.$
Combining the three inequalities above and utilizing the assumption $\eta\leq 1/L$, we complete the proof. ∎
6.1 Hidden Convex Setting
Combining theorem 7 with lemma 2, we obtain the following theorem.
Theorem 8.
[…]
Proof.
By Theorem 7, subtracting $F(x^{*})$ from both sides of (18), setting $\mu_{H}=0$, and taking the expectation, we have for any $\eta\leq 1/L$ that
${\mathbb{E}}\left[F(x^{t+1})-F(x^{*})\right]$$\leq$ $(1-\alpha){\mathbb{E}}\left[F(x^{t})-F(x^{*})\right]+\frac{\alpha^{2}}{\mu_{c}^{2}\eta}{\mathbb{E}}\left[\left\|c(x^{t})-c(x^{*})\right\|^{2}\right]$
$\qquad-\left(\frac{1}{2\eta}-\frac{L}{2}\right){\mathbb{E}}\left[\left\|x^{t+1}-x^{t}\right\|^{2}\right]+\frac{\eta}{2}{\mathbb{E}}\left[\left\|g^{t}-\nabla F(x^{t})\right\|^{2}\right]$
$\leq$ $(1-\alpha){\mathbb{E}}\left[F(x^{t})-F(x^{*})\right]+\frac{\alpha^{2}D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\eta}$
$\qquad-\left(\frac{1}{2\eta}-\frac{L}{2}\right){\mathbb{E}}\left[\left\|x^{t+1}-x^{t}\right\|^{2}\right]+\frac{\eta}{2}{\mathbb{E}}\left[\left\|g^{t}-\nabla F(x^{t})\right\|^{2}\right],$
where the second inequality uses boundedness of $\mathcal{U}$.
Summing up the inequality above with a $\frac{\eta}{\beta}$ multiple of the result of Lemma 2, we recognize the Lyapunov function $\Lambda_{t}^{\text{HB}}$ defined in (17), and derive
$\Lambda_{t+1}^{\text{HB}}$$\leq$ $\Lambda_{t}^{\text{HB}}-\alpha{\mathbb{E}}\left[F(x^{t})-F(x^{*})\right]-\frac{\eta}{2}{\mathbb{E}}\left[\left\|g^{t}-\nabla F(x^{t})\right\|^{2}\right]$
$\qquad+\frac{\alpha^{2}D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\eta}-\left(\frac{1}{2\eta}-\frac{L}{2}-\frac{3L^{2}\eta}{\beta^{2}}\right){\mathbb{E}}\left[\left\|x^{t+1}-x^{t}\right\|^{2}\right]+\beta\eta\sigma^{2}$
$\leq$ $(1-\alpha)\Lambda_{t}^{\text{HB}}+\frac{\alpha^{2}D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\eta}+\beta\eta\sigma^{2},$
where the last step holds for $\alpha\leq\beta/2$ and $\eta\leq\frac{\beta}{4L}$. Unrolling the recursion from $t=0$ to $t=T-1$ and choosing $\eta=\frac{\beta}{4L}$, we obtain
$\Lambda_{T}^{\text{HB}}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}^{\text{HB}}+\frac{\alpha D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\eta}+\frac{\beta\eta\sigma^{2}}{\alpha}$
$\leq$ $(1-\alpha)^{T}\Lambda_{0}^{\text{HB}}+\frac{4LD_{\mathcal{U}}^{2}}{\mu_{c}^{2}}\frac{\alpha}{\beta}+\frac{\sigma^{2}}{4L}\frac{\beta^{2}}{\alpha}\leq\varepsilon,$
where the last inequality holds by setting $\alpha=\min\left\{\frac{\beta}{2},\frac{3\mu_{c}^{2}}{2LD_{\mathcal{U}}^{2}}\beta\varepsilon,\frac{\sigma\mu_{c}}{4LD_{\mathcal{U}}}\beta^{\frac{3}{2}}\right\}$ , $\beta=\min\left\{1,\frac{\mu_{c}^{2}}{9D_{\mathcal{U}}^{2}\sigma^{2}}\varepsilon^{2}\right\}$, and the number of iterations as
$T=\frac{1}{\alpha}\log\left(\frac{3\Lambda_{0}^{\text{HB}}}{\varepsilon}\right)=\widetilde{\mathcal{O}}\left(\frac{LD_{\mathcal{U}}^{2}}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{LD_{\mathcal{U}}^{4}\sigma^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right).$
6.2 Hidden Strongly Convex Setting
We conclude the section with the improved result for Projected SGD with momentum under hidden strong convexity.
Theorem 9.
[…]
Proof.
Applying Theorem 7 with $\mu_{H}>0$, and setting $\alpha$ small enough allows us to cancel the term invloving $\left\|c(x^{t})-c(x^{*})\right\|^{2}$. The rest of the proof is similar to the one of Theorem 6. ∎
We remark that both Theorems 8 and 9 provide last iterate global convergence for Projected SGD with momentum without the need of using large mini-batch. Additionally, the gradient estimate $g^{t}$ is guaranteed to converge to the true gradient $\nabla F(x^{*})$ at the optimum $x^{*}\in\mathcal{X}^{*}$, which might be non-zero when minimizing over a compact set $\mathcal{X}$. In the hidden strongly convex case, similarly to Corollaries 2 and 3, the result of Theorem 9 can be translated to the point convergence to the optimal solution.
7 Conclusions
In this work, we study stochastic optimization under hidden convexity and develop sample complexity results for batch-free stochastic (sub-) gradient methods with projection.
Several questions remain open. 1) We know that in case $\mu_{H}>0$, the derived sample complexity is worst-case optimal (up to the logarithmic factor) in terms of dependence on $\varepsilon$ since it matches the optimal rate known for strongly convex $F(\cdot)$, and therefore, the complexity bounds are unimprovable for SM and P-SGD. However, for merely convex $H(\cdot)$, i.e., $\mu_{H}=0$, it is unclear if our $\widetilde{\mathcal{O}}(\varepsilon^{-3})$ sample complexity is tight for SM and P-SGD. 2). The benefits of momentum variants of P-SGD can be further explored, e.g., to understand if Nesterov’s acceleration is possible under hidden convexity. 3) When $\mu_{H}=0$, our iteration and sample complexity results depend on the diameter of the reformulated problem. It would be interesting to explore if $D_{\mathcal{U}}$ can be replaced with the distance to the solution, i.e., $\left\|c(x^{0})-c(x^{*})\right\|$.
There are also many other directions to explore in the future. 1) SM and P-SGD are the simplest and generic methods for solving (1). It is important to explore more advanced specialized algorithms for applications, which may potentially speed up the convergence. For instance, given a stochastic information about the map $c(x)={\mathbb{E}}\left[c(x,\xi)\right]$, one can utilize the samples $c(x,\xi)$ or $\nabla c(x,\xi)$ in the algorithm. Despite some recent progress [15], the rigorous validation of such methods remains an open problem with a general convex constrained $\mathcal{X}$. 2) The development of stochastic gradient methods for solving hidden convex problems with non-convex (e.g., hidden convex) constraints is an interesting research direction [86, 25, 96, 53]. 3) Extension of our results to hidden convex saddle point problems and games [68] also merits further exploration.
Appendix A Technical Lemma
We report the following technical lemma from [22, 34] and include their slightly modified proofs for completeness.
Lemma 3 (Lemma 3.2 in [22]).
Let $\rho>\ell$, and for any $x^{t}\in\mathcal{X}$, define $\hat{x}^{t}:=\operatorname{prox}_{\Phi/\rho}(x^{t})$, where $\Phi:=F+\delta_{\mathcal{X}}$. Then $\hat{x}^{t}=\Pi_{\mathcal{X}}\left(\eta\rho x^{t}-\eta\hat{g}^{t}+(1-\eta\rho)\hat{x}^{t}\right),$ where $\hat{g}^{t}\in\partial F(\hat{x}^{t})$.
Proof.
By definition of $\hat{x}^{t}$ and $\Phi(\cdot)$, we have
$0\in\partial\left(F+\frac{\rho}{2}\left\|\cdot-x^{t}\right\|^{2}+\delta_{\mathcal{X}}\right)(\hat{x}^{t})=\hat{g}^{t}+\rho\left(\hat{x}^{t}-x^{t}\right)+\partial\delta_{\mathcal{X}}\left(\hat{x}^{t}\right),$
where the last equality holds, since $F(\cdot)+\frac{\rho}{2}\left\|\cdot-x^{t}\right\|^{2}$, and $\delta_{\mathcal{X}}(\cdot)$ are both convex (due to the conic combination rule). Multiplying both sides by $\eta>0$ and rearranging, we get $z^{t}:=\eta\rho x^{t}-\eta\hat{g}^{t}+(1-\eta\rho)\hat{x}^{t}\in\hat{x}^{t}+\eta\partial\delta_{\mathcal{X}}\left(\hat{x}^{t}\right).$ Therefore, by the optimality condition for the proximal sub-problem, we have $\hat{x}^{t}=\operatorname{prox}_{\eta\delta_{\mathcal{X}}}\left(z^{t}\right)=\Pi_{\mathcal{X}}(z^{t})$. ∎
Proof of Lemma 1. Lemma 3 states that for any $\hat{g}^{t}\in\partial F(\hat{x}^{t})$ and $z^{t}=\eta\rho x^{t}-\eta\hat{g}^{t}+(1-\eta\rho)\hat{x}^{t}$, we have $\hat{x}^{t}=\Pi_{\mathcal{X}}(z^{t})$. Thus, using the update rule of $x^{t+1}$ and non-expansiveness of the projection, we derive
$\mathbb{E}\left[\left\|x^{t+1}-\hat{x}^{t}\right\|^{2}\mid x^{t}\right]=\mathbb{E}\left[\left\|\Pi_{\mathcal{X}}\left(x^{t}-\eta g(x^{t},\xi^{t})\right)-\Pi_{\mathcal{X}}\left(z^{t}\right)\right\|^{2}\mid x^{t}\right]$
$\leq\mathbb{E}\left[\left\|x^{t}-\eta g(x^{t},\xi^{t})-\left(\eta\rho x^{t}-\eta\hat{g}^{t}+(1-\eta\rho)\hat{x}^{t}\right)\right\|^{2}\mid x^{t}\right]$
$=\mathbb{E}\left[\left\|(1-\eta\rho)\left(x^{t}-\hat{x}^{t}\right)-\eta\left(g(x^{t},\xi^{t})-\hat{g}^{t}\right)\right\|^{2}\mid x^{t}\right]$
=$\overset{(i)}{=}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-(1-\eta\rho)\eta\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle+\eta^{2}{\mathbb{E}}\left[\left\|g(x^{t},\xi^{t})-\hat{g}^{t}\right\|^{2}\mid x^{t}\right]$
$\overset{(ii)}{\leq}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-(1-\eta\rho)\eta\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle+2G_{F}^{2}\eta^{2}$
$\overset{(iii)}{\leq}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-(1-\eta\rho)\eta\ell\left\|x^{t}-\hat{x}^{t}\right\|^{2}+2G_{F}^{2}\eta^{2}$
$\leq(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+\eta^{2}G_{F}^{2},$
where in $(i)$ we use unbiasedness of the gradient estimator. In $(ii)$, we use Young’s inequality and A.2, $(iii)$ holds by hypo-monotonicity inequality $\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle\geq-\ell\left\|x^{t}-\hat{x}^{t}\right\|^{2}$. The last inequality holds by the choice of $\rho$ and $\eta$square
Lemma 4 (Lemma 3.4 in [22]).
Let A.1’, A.2’ hold, and $\rho=4L$, $\eta\leq\frac{2}{9L}$. Then for all $t\geq 0$
$\mathbb{E}\left[\left\|x^{t+1}-\hat{x}^{t}\right\|^{2}\mid x^{t}\right]\leq(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+\sigma^{2}\eta^{2}$
Proof.
For a differentiable $F(\cdot)$ Lemma 3 implies that for $z^{t}=\eta\rho x^{t}-\eta\nabla F\left(\hat{x}^{t}\right)+(1-\eta\rho)\hat{x}^{t}$, we have $\hat{x}^{t}=\Pi_{\mathcal{X}}(z^{t})$. Thus, using the update rule of $x^{t+1}$ and non-expansiveness of the projection, we derive
$\mathbb{E}\left[\left\|x^{t+1}-\hat{x}^{t}\right\|^{2}\mid x^{t}\right]=\mathbb{E}\left[\left\|\Pi_{\mathcal{X}}\left(x^{t}-\eta\nabla f\left(x^{t},\xi^{t}\right)\right)-\Pi_{\mathcal{X}}\left(z^{t}\right)\right\|^{2}\mid x^{t}\right]$
$\leq\mathbb{E}\left[\left\|x^{t}-\eta\nabla f\left(x^{t},\xi^{t}\right)-\left(\eta\rho x^{t}-\eta\nabla F\left(\hat{x}^{t}\right)+(1-\eta\rho)\hat{x}^{t}\right)\right\|^{2}\mid x^{t}\right]$
$=\mathbb{E}\left[\left\|(1-\eta\rho)\left(x^{t}-\hat{x}^{t}\right)-\eta\left(\nabla f\left(x^{t},\xi^{t}\right)-\nabla F\left(\hat{x}^{t}\right)\right)\right\|^{2}\mid x^{t}\right]$
$=\mathbb{E}\left[\left\|(1-\eta\rho)\left(x^{t}-\hat{x}^{t}\right)-\eta\left(\nabla F\left(x^{t}\right)-\nabla F\left(\hat{x}^{t}\right)\right)-\eta\left(\nabla f\left(x^{t},\xi^{t}\right)-\nabla F\left(x^{t}\right)\right)\right\|^{2}\mid x^{t}\right]$
=$\overset{(i)}{=}\left\|(1-\eta\rho)\left(x^{t}-\hat{x}^{t}\right)-\eta\left(\nabla F\left(x^{t}\right)-\nabla F\left(\hat{x}^{t}\right)\right)\right\|^{2}+\eta^{2}\mathbb{E}\left[\left\|\nabla f\left(x^{t},\xi^{t}\right)-\nabla F\left(x^{t}\right)\right\|^{2}\mid x^{t}\right]$
$\overset{(ii)}{\leq}\left\|(1-\eta\rho)\left(x^{t}-\hat{x}^{t}\right)-\eta\left(\nabla F\left(x^{t}\right)-\nabla F\left(\hat{x}^{t}\right)\right)\right\|^{2}+\eta^{2}\sigma^{2}$
$=(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-2(1-\eta\rho)\eta\left(x^{t}-\hat{x}^{t},\nabla F\left(x^{t}\right)-\nabla F\left(\hat{x}^{t}\right)\right\rangle+\eta^{2}\sigma^{2}$
$\quad+\eta^{2}\left\|\nabla F\left(x^{t}\right)-\nabla F\left(\hat{x}^{t}\right)\right\|^{2}$
$\overset{(iii)}{\leq}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}+2(1-\eta\rho)\eta L\left\|x^{t}-\hat{x}^{t}\right\|^{2}+\eta^{2}L^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}+\eta^{2}\sigma^{2}$
$=(1-\eta\rho)\left(1-\eta\rho+2\eta L+\frac{\eta^{2}L^{2}}{1-\eta\rho}\right)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+\eta^{2}\sigma^{2}$
$\leq(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+\eta^{2}\sigma^{2},$
where in $(i)$ and $(ii)$ use unbiasedness of the gradient estimator and bounded variance. In $(iii)$, we use Cauchy–Schwarz inequality and smoothness of $F(\cdot)$, i.e., $\left\|\nabla F\left(\hat{x}^{t}\right)-\nabla F\left(x^{t}\right)\right\|\leq L\left\|\hat{x}^{t}-x^{t}\right\|$. The last inequality holds by the choice of $\rho,\eta$ and $2\eta L\leq\frac{\eta\rho}{2}$, and $\frac{\eta^{2}L}{1-\eta\rho}\leq\frac{\eta\rho}{2}$. ∎
Proof of Lemma 2. Using the update rule of $g^{t+1}$ and the unbiasedness of stochastic gradients, we have
$\mathbb{E}$ $\left[\left\|g^{t+1}-\nabla F(x^{t+1})\right\|^{2}\right]={\mathbb{E}}\left[\left\|(1-\beta)g^{t}+\beta\nabla f(x^{t+1},\xi^{t+1})-\nabla F(x^{t+1})\right\|^{2}\right]$
$=(1-\beta)^{2}{\mathbb{E}}\left[\left\|g^{t}-\nabla F(x^{t+1})\right\|^{2}\right]+\beta^{2}{\mathbb{E}}\left[\left\|\nabla f(x^{t+1},\xi^{t+1})-\nabla F(x^{t+1})\right\|^{2}\right]$
$\leq(1-\beta)^{2}\left(1+\nicefrac{{\beta}}{{2}}\right){\mathbb{E}}\left[\left\|g^{t}-\nabla F(x^{t})\right\|^{2}\right]+\left(1+\nicefrac{{2}}{{\beta}}\right){\mathbb{E}}\left[\left\|\nabla F(x^{t})-\nabla F(x^{t+1})\right\|^{2}\right]+\beta^{2}\sigma^{2}$
$\leq(1-\beta){\mathbb{E}}\left[\left\|g^{t}-\nabla F(x^{t})\right\|^{2}\right]+\frac{3L^{2}}{\beta}{\mathbb{E}}\left[\left\|x^{t}-x^{t+1}\right\|^{2}\right]+\beta^{2}\sigma^{2},$
where the first inequality uses Young’s inequality and the bound of the variance of stochastic gradients, and the last step uses the Lipschitz continuity of the gradient and the fact that $(1-\beta)\left(1+\nicefrac{{\beta}}{{2}}\right)\leq 1$ for all $\beta\in(0,1]$. $\square$
The following technical lemma is fairly standard, e.g., see [41].
Lemma 5.
Let ϕ()italic-$\phi(\cdot)$ be convex and for some $\eta>0$, $x\in\mathcal{X}$, define $x^{+}:=\operatorname*{arg\,min}_{y\in\mathcal{X}}\left\{\phi(y)+\frac{1}{2\eta}\left\|y-x\right\|^{2}\right\}$, then
$\phi(y)+\frac{1}{2\eta}\left\|y-x\right\|^{2}\geq\phi(x^{+})+\frac{1}{2\eta}\left\|x^{+}-x\right\|^{2}+\frac{1}{2\eta}\left\|y-x^{+}\right\|^{2}\quad\text{for all }y\in\mathcal{X}.$
Appendix B Further Improvements with Mini-batching
Using only one (post-processing) step of mini-batch P-SGD to the output of one batch P-SGD is sufficient to translate convergence from $\Phi_{1/4L}(\cdot)$ to $F(\cdot)$.
Corollary 4.
Let the assumptions of Theorem 5 hold and $G_{F}>0$ be the Lipschitz constant of $F(\cdot)$ over $\mathcal{X}$. Set $x^{T+1}=\Pi_{\mathcal{X}}\left(x^{T}-\frac{1}{3L}\frac{1}{B_{0}}\sum_{i=1}^{B_{0}}\nabla F(x^{T},\xi_{i}^{T})\right)$, where $B_{0}\geq\min\{1,\left(\frac{G_{F}\sigma}{3L\varepsilon}\right)^{2}\}$, $x^{T}$ is the output of (15) applied with batch-size $B=1$ after $T$ iterations (given by Theorem 5). Then ${\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq 2\,\varepsilon$.
Proof.
Define $x_{+}^{T}:=\Pi_{\mathcal{X}}(x^{T}-\frac{1}{\rho-L}\nabla F(x^{T}))$, $\rho=4L$. Notice that $F(x_{+}^{T})=\Phi(x_{+}^{T})\leq\Phi_{1/\rho}(x^{T})$, where the inequality follows by [77, Poposition 2.5-(i)] with $\gamma:=(\rho-L)^{-1}$. Therefore, Theorem 5 implies that
${\mathbb{E}}\left[F(x_{+}^{T})-F(x^{*})\right]\leq{\mathbb{E}}\left[\Phi_{1/\rho}(x^{T})-F(x^{*})\right]\leq\varepsilon,$
On the other hand, the post-processing step guarantees
$\mathbb{E}$ $\left[F(x^{T+1})-F(x_{+}^{T})\right]\leq G_{F}\,{\mathbb{E}}\left[\left\|x^{T+1}-x_{+}^{T}\right\|\right]$
$\leq G_{F}{\mathbb{E}}\left[\left\|\Pi_{\mathcal{X}}\left(x^{T}-\frac{1}{3L}\frac{1}{B_{0}}\sum_{i=1}^{B_{0}}\nabla F(x^{T},\xi_{i}^{T})\right)-\Pi_{\mathcal{X}}\left(x^{T}-\frac{1}{3L}\nabla F(x^{T})\right)\right\|\right]$
$\leq\frac{G_{F}}{3L}{\mathbb{E}}\left[\left\|\frac{1}{B_{0}}\sum_{i=1}^{B_{0}}\nabla F(x^{T},\xi_{i}^{T})-\nabla F(x^{T})\right\|\right]\leq\frac{G_{F}\sigma}{3L\sqrt{B_{0}}}\leq\varepsilon.$
Combining the above two inequalities, the result follows. ∎
The following corollary shows that if we apply mini-batching at each iteration with sufficiently large batch-size, then the number of iterations required for convergence is reduced to $\widetilde{\mathcal{O}}(\varepsilon^{-1})$.
Corollary 5.
Let the assumptions of Theorem 5 hold and $G_{F}>0$ be the Lipschitz constant of $F(\cdot)$ over $\mathcal{X}$. Suppose P-SGD with batch-size $B$ is applied, i.e., $\{x^{t}\}_{t\geq 0}$ is generated by $x^{t+1}=\Pi_{\mathcal{X}}\left(x^{t}-\eta\frac{1}{B}\sum_{i=1}^{B}\nabla F(x^{t},\xi_{i}^{t})\right)$ with $\eta=\frac{2}{9L}$, $B\geq\min\{1,\frac{D_{\mathcal{U}}^{2}\sigma^{2}}{\mu_{c}^{2}\varepsilon^{2}}\}$. Define $x^{T+1}=\Pi_{\mathcal{X}}\left(x^{T}-\frac{1}{3L}\frac{1}{B_{0}}\sum_{i=1}^{B_{0}}\nabla F(x^{T},\xi_{i}^{T})\right)$, where $\rho=4L$, and $B_{0}\geq\min\{1,\left(\frac{G_{F}\sigma}{3L\varepsilon}\right)^{2}\}$. Then ${\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq 2\,\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{LD_{\mathcal{U}}^{2}}{\mu_{c}^{2}\varepsilon}\right).$
Proof.
The proof follows from the previous corollary by replacing $\sigma^{2}$ with $\sigma^{2}/B$. ∎
However, we highlight that the results of theorems 5 and 4 do not require using large batches of samples at every iteration.
Appendix C Historical Remarks
The sub-gradient method, its special case, Projected SGD, and their numerous variants have a long history of development since the first works on stochastic approximation appeared in 1950s [73, 49, 9, 18]. Convex optimization. The case of convex $F(\cdot)$ is particularly well documented [63, 1, 37]. Researchers have studied how to deal with convex constraints, proximal operators, general Bregman divergences [64, 5], and leveraging averaging and momentum schemes [70, 76, 38, 55]. In the convex case, the global convergence of gradient methods in the function value, i.e., find $x\in\mathcal{X}$ with ${\mathbb{E}}\left[F(x)-F(x^{*})\right]\leq\varepsilon$ for any $\varepsilon>0$, is naturally possible and the sample complexity required is ${\mathcal{O}}\left(\varepsilon^{-2}\right)$. [^1] For P-SGD under smooth and bounded variance assumptions, A.1’ and A.2’ in section 2, or for SM under Lipschitz continuity and bounded second moment of stochastic sub-gradients, i.e., A.2.
Non-convex optimization. In the last decade, the interest in the optimization community shifted towards general non-convex problems (often smooth or weakly convex), where only convergence to a FOSP is possible in general [48, 27, 3, 87], i.e., find $x\in\mathcal{X}$ with ${\mathbb{E}}\left[\left\|\nabla F(x)\right\|\right]\leq\epsilon$ when $F(\cdot)$ is smooth. Similar to developments in convex optimization, convergence of non-convex SGD extends to constrained/proximal setting [39, 51, 11], mirror descent [95, 23], momentum [57, 34], variance-reduction [21, 3], and biased gradient setting [45, 43, 44]. For the more general weakly-convex case [22, 60, 95], the convergence guarantees are usually with respect to a gradient norm of a smoothed objective. Some works consider non-convex functions with a specific compositional structure similar to (2), e.g. the composition of a convex function with a differentiable and smooth map $c(\cdot)$, see [65, 52, 28, 94]. Recently a number of works focus on non-convex non-smooth optimization (beyond weak convexity) and develop convergence for suitably defined notions of FOSP [20, 91, 46]. Although the above works consider non-convex problems, which find a wide range of applications, they often only provide convergence to a FOSP rather than global convergence in the function value.